OpenStudy (anonymous):

how do i find the following : (f+h)(1) if f(x)={(3,5), (2,4), (1,7)} h(x)={(3,2), (4,3), ( 1, 6)}

OpenStudy (anonymous):

Since you know what the answer of f(1) is and h(1), because they are given to you in the sets, you just add the results together.

OpenStudy (anonymous):

6+7=13

myininaya (myininaya):

i agree

myininaya (myininaya):

just like (f+h)(3)=f(3)+h(3)=5+2+7

myininaya (myininaya):

that last plus sign is suppose to be an equal sign

myininaya (myininaya):

5+2=7*